OpenStudy (dinmosiaren):

@cody_123 @kagıtucak Thirty is 12% of what number? A. 2.5 B. 360 C. 3.6 D. 250

OpenStudy (anonymous):

12/100 *x =30 x=30*100 /12=250

OpenStudy (zzr0ck3r):

\(30 = .12x\) solve x

OpenStudy (dinmosiaren):

\[WHAT?\]

OpenStudy (anonymous):

\[\frac{ 12 }{ 100}*x=30\]

OpenStudy (anonymous):

x=250

OpenStudy (dinmosiaren):

Thanks